# UC-MOA: Utility-Conditioned Multi-Objective Alignment for Distributional Pareto-Optimality

### v1

@@ 3.2 Utility Index Selection during Training
To robustly determine the optimal utility index during training, we employ a percentile ranking strategy in both the offline and online phases. For every sample with reward vector $\mathbf{z}$, each candidate utility function $\hat{g}_{\psi_{i}}$ produces a utility score. We then compute the percentile rank for the $i$-th utility function as
@@
$\text{percentile}_{i}=\frac{\text{rank of }\hat{g}_{\psi_{i}}(\mathbf{z})}{N},$ (5)
where $N$ is the total number of samples, and the rank is determined by the position of $\hat{g}_{\psi_{i}}(\mathbf{z})$ in the sorted array of scores for that function. The percentile ranking provides a normalized measure of performance for each utility function, mitigating the effects of scale differences and ensuring that selection is robust across diverse reward distributions. During offline and online training, the utility index is then selected as the one with the highest percentile rank. We present a concrete example in Appendix B.2.
3.3 Offline Training
In the offline phase, our approach establishes a robust grounding of the language model in distributional reward signals by embedding an explicit symbolic indicator of the preferred utility function directly into the training prompt. For each training sample, denoted by $(x,y)$ with $x$ representing the input prompt and $y$ the corresponding response, we first compute a multi-objective return vector $\mathbf{z}(x,y)$ and then evaluate a suite of strictly increasing candidate utility functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ on $\mathbf{z}(x,y)$. The scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z}(x,y))$ for $i=1,\dots,M$ are then obtained, and for each function, we calculate its percentile rank over the training distribution. The maximum utility index is determined as the index with the highest percentile:
$i^{*}=\arg\max_{i\in\{1,\dots,M\}}\left\{\text{percentile}_{i}\right\}.$ (6)
We augment the original prompt $x$ by appending a specialized token that encodes this index; that is, the modified prompt is given by
$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*}.$
The model is then fine-tuned via a standard cross-entropy loss:
@@ 3.3 Offline Training
$\mathcal{L}_{\text{offline}}(\theta)=-\sum_{t}\log\pi_{\theta}(y_{t}\mid x^{\prime},y_{<t}),$
ensuring that the model learns to internalize the mapping between input contexts and their corresponding optimal distributional reward profiles. This training stage leverages supervised techniques exclusively, thereby avoiding the instability often associated with RL-based methods.
3.4 Online Fine-Tuning
To enhance the generalizability of the fine-tuned model, we incorporate additional online stages. In the online stage, the large language model generates additional candidate responses and we selectively choose those responses that exhibit superior distributional reward characteristics to fine-tune the model. Given an input prompt $x$, the model produces a response $y$ whose multi-objective return $\mathbf{z}(y)$ is computed and then evaluated by each candidate utility function, yielding scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z}(y))$ for $i=1,\dots,M$. For each candidate response, we compute the percentile rank of each utility score as described above, and the optimal utility index is selected as the one with the highest percentile as in Equation (5), resulting in an updated input
$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*}.$
The response $y$ and its corresponding modified prompt $x^{\prime}$ are stored in an online training buffer for further model updates. Our online training process also incorporates a rejection sampling mechanism to filter out responses that do not exceed predetermined utility thresholds across multiple reward dimensions. The selected samples are then used to update the language model via the same reward-conditional supervised objective as in the offline training phase.
@@ 3.5 Inference Stage
At inference time, the user specifies a preference vector $\mathbf{w}=[w_{1},\ldots,w_{N}]$ that encodes the relative importance of each reward dimension $r_{1},\ldots,r_{N}$. To translate these preferences into a concrete reward profile, we apply a linear preference-to-reward mapping for each dimension:
$f_{i}(w_{i})=w_{i}\times(r_{i}^{\mathrm{max}}-r_{i}^{\mathrm{min}})+r_{i}^{\mathrm{min}},$
where $r_{i}^{\mathrm{max}}$ and $r_{i}^{\mathrm{min}}$ denote the maximum and minimum values of the $i$-th reward observed in the training dataset. This mapping yields a target reward vector $\mathbf{z}=[f_{1}(w_{1}),\dots,f_{N}(w_{N})]$ that accurately reflects the user’s preferences while normalizing the reward scales across dimensions.
Subsequently, the target reward vector $\mathbf{z}$ is input into each of the pre-trained, strictly increasing utility functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ to obtain scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z})$ for $i=1,\dots,M$. at inference time the reward distribution is more stable and the set of candidate responses is smaller; therefore, we directly use the argmax operation to select the index corresponding to the maximum expected utility:
$i^{*}=\arg\max_{i\in\{1,\dots,M\}}U_{i}.$ (7)
This index $i^{*}$ is encoded into the input prompt via a symbolic token—specifically, the prompt is augmented as
$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*},$
thereby conditioning the language model to generate responses that are robustly aligned with the user’s specified multi-objective reward preferences while obviating the need for direct numerical processing.
@@ 4 Evaluation
In this section, we start with the experimental setup, including the tasks, baselines, and implementation. We then present our experiment design, followed by experiment results and analysis. We provide our code in the anonymous link https://anonymous.4open.science/r/UC-MOA-8D36.
@@ 4.1 Experiment Setup
Tasks. Following RiC (Yang et al., 2024c), we select two text generation tasks for our experiments, i.e., Helpful Assistant task (Bai et al., 2022) and the Reddit Summary task (Stiennon et al., 2020). For the Helpful Assistant task, we utilize three open-sourced reward models on Huggingface to measure three different objectives “harmless”, “helpful”, and “humor”. We focus on three multi-objective subtasks: (1) “harmless” vs “helpful”; (2) “humor” vs “helpful”; (3) “harmless” vs “humor”. Regarding the Reddit Summary task, we consider three reward models namely “summary”, “hallucinated”, and “faithful” to evaluate different perspectives of the summary quality. We focus on three multi-objective subtasks: (1) “summary” vs “faithful”; (2) “hallucinated” vs “faithful”; (3) “summary” vs “hallucinated”. We provide more details about the datasets and the reward models in Appendix A.
@@
Baseline methods.We compare our method with two state-of-the-art multi-objective RLHF methods, Rewarded Soups, and Reward-in-Context (RiC). We also report the performance of the base model and the supervised fine-tuned (SFT) model.
@@
Implementation. We implement the proposed method using PyTorch. We train the machine learning models on a server with 8 NVIDIA A100 80GB GPUs and 4TB memory for all methods. We select the LLaMA 2 7B-hf model as the base model for alignment following the setting of Yang et al. (2024c). We list the implementation details (e.g., network structure for the parameterized utility function, hyper-parameter choices) in Appendix B.
4.2 Experiment Design
We use the following experiments to evaluate the effectiveness and efficiency of our method, and other factors that influence the performance (e.g., alternative design choices, hyper-parameters).
Experiment I. To evaluate the effectiveness of our method, we compare the curves of the multi-dimensional average test rewards that correspond to the empirical Pareto fronts generated by different methods. A method demonstrating an outer curve indicates better performance, as it captures a wider range of trade-offs across objectives. We plot the Pareto fronts of all baseline methods across the Helpful Assistant task and the Reddit Summary task.
Experiment II. We analyze the computational efficiency of each method by recording the training time of each baseline method. We run experiments with NVIDIA A100 80GB GPUs and report the GPU hours utilized by each method.
Experiment III. To validate the necessity of the online stage, we perform an ablation study to compare the performance of pure offline training (“w/o online”), a single online iteration (“w/ online iter 1”), and two online iterations (“w/ online iter 2”).
Experiment IV. We assess the consistency of our method during the inference phase by measuring the alignment between the user-specified target objective and the reward outcomes obtained from the LLM’s generated responses. Each prompt provided to the LLM embeds a target objective requirement (e.g., in the form <max_utility_index> a, where the letter encodes the desired objective). During evaluation, the LLM is tasked with generating responses that adhere to these preset requirements. For each generated response, we compute a set of utility values using a suite of ten reward models. To account for scale differences across objectives, the computed reward values are first normalized on a per-utility basis. We then quantify consistency by comparing the expected utility and the value distribution of each utility function with the target objective extracted from the prompt.
Experiment V. To investigate how the number of utility functions impacts the performance of our method, we vary the number of utility functions from {5,10,15} and test the performance of our method with different numbers of utility functions.
Experiment VI. We investigate an alternative design choice of the utility functions, i.e., linear functions. The linear function is designed in the form of $w^{T}x$ where $||w||_{2}=1$. To construct $N$ linear utility functions, we uniformly sample $N$ points from the unit circle and set $w$ accordingly. We plot the Pareto fronts using linear utility functions and our method across all tasks.
Experiment VII. We also investigate two metrics developed specifically for assessing distributional preference, i.e., Constraint Satisfaction and Variance Objective as in previous distributional studies (Cai et al., 2023). They are designed to underscore the optimality of multivariate distributions associated with LLMs. In terms of Constraint Satisfaction, we randomly generate $M=100$ constraints. The Constraint Satisfaction metric is then computed by considering the highest probability within the policy set that satisfies each individual constraint. The Variance Objective metric, on the other hand, involves generating $M=100$ random linear weights. These weights are applied to both the expected returns and the standard deviations of returns in each dimension. This objective encourages attaining greater expected returns while simultaneously increasing variance, thereby catering to dynamic preferences. Further details about the implementation of these evaluation metrics are provided in Appendix B.4.
4.3 Experiment Results.
Performance comparison. We compare the multi-dimensional average test rewards obtained from the empirical Pareto fronts generated by different methods on two distinct tasks: the Helpful Assistant task and the Reddit Summary task. We show the results of “harmless” vs. “humor” task and “deberta” vs. “faithful” in Figure 2 while leaving the rest of the results in Appendix C.1. We observe that our method consistently produces an outer Pareto curve that dominates those of Rewarded Soups and RiC, indicating its superior ability to capture a broader spectrum of trade-offs.
Figure 2: Pareto fronts of “harmless” vs “humor” and “deberta” vs. “faithful” tasks. Each marker on the plotted curves represents the average test rewards corresponding to a particular user preference.
Computational cost. We measure the time complexity of all methods in the two-objective Helpful Assistant task and the two-objective Reddit Summary task using NVIDIA A100 80GB GPUs. We report the time cost (in the form of GPU hours) in Table 1 and Table 2. We observe that Rewarded Soups consistently consumes the most GPU hours across both tasks, while our method (w/ online iter2) achieves notably lower costs. Meanwhile, RiC requires more computing time than our method, though it remains more cost-friendly than Rewarded Soups. These findings demonstrate that our approach not only delivers strong performance (as shown in previous experiments) but also offers a clear advantage in terms of computational efficiency.
  Method GPU hours
  Rewarded Soups 69.5
RiC 20.3
Ours w/o online 8.9
Ours w/ online iter1 10.1
Ours w/ online iter2 11.3
Table 1: Comparison of GPU hours in the Helpful Assistant task, where the number of preference $M=10$, number of rewards $N=2$.
  Method GPU hours
  Rewarded Soups 15.9
RiC 5.7
Ours w/o online 2.0
Ours w/ online iter1 2.4
Ours w/ online iter2 2.8
Table 2: Comparison of GPU hours in the Reddit Summary task, where the number of preference $M=10$, number of rewards $N=2$.
Necessity of online stage. We evaluate the impact of the online stage by comparing models trained purely offline (“w/o online”) with those undergoing one (“w/ online iter 1”) and two online iterations (“w/ online iter 2”) across all tasks. We present the performance results for “harmless” vs. “humor” and “summary” vs. “faithful” tasks in Figure 3 while leaving the rest of the results in Appendix C.2. We observe that models incorporating online iterations consistently outperform the offline-only baseline. Notably, even a single online iteration yields significant improvements, with further performance gains observed when two iterations are applied, thereby underscoring the necessity of the online stage in our training framework. The online stage is important because it enables the model to generalize over unseen samples in the offline dataset, and ultimately enhancing its ability to generate appropriate responses corresponding to the required utility function index.
Figure 3: Ablation study of the online stage on “harmless” vs. “humor” task and “summary” vs. “faithful” task.

### v2

@@ 3.2 Utility Index Selection via Percentile Ranking
To determine which pre-defined personalized preference profile best matches a given response’s reward characteristics during training, we employ a percentile ranking strategy based on utility functions. For each sample with reward vector $\mathbf{z}$, we compute the utility score $u_{i}=\hat{g}_{\psi_{i}}(\mathbf{z})$ for each function $i$. The percentile rank for the $i$-th utility function as
@@
$\text{percentile}j(\mathbf{z})=\frac{\text{rank of }\hat{g}_{\psi_{j}}(\mathbf{z})\text{ w.r.t. its scores on dataset }\mathcal{D}_{\text{ref}}}{|\mathcal{D}_{\text{ref}}|},$ (5)
where $\mathcal{D}_{\text{ref}}$ is a reference dataset, such as the offline training set or the current online buffer. Similar percentile-based shaping is a standard trick in which it makes the update rule invariant to any monotone transformation of the objective and empirically stabilizes learning under widely varying or drifting reward scales [22, 23]. The selected utility index $i^{*}$ corresponds to the utility function that yields the highest percentile rank. An illustrative example is provided in Appendix D.2.
3.3 Offline Training: Learning to Map Symbolic Tokens to Personalized Styles
In the offline phase, we leverage a fixed dataset $\mathcal{D}$ of prompt-response pairs $(x,y)$ to teach the LLM to associate symbolic utility tokens with specific personalized response styles. For each training sample $(x,y)$, we compute its normalized reward vector $\mathbf{z}(x,y)\in[0,1]^{K}$. Using percentile ranking (Equation (5) relative to $\mathcal{D}$), we identify the index $i^{*}$ of the utility function $g_{\psi_{i^{*}}}$ for which $y$ achieves the highest relative performance, indicating that $y$ best embodies the preference structure $i^{*}$.
The prompt $x$ is augmented with a symbolic token representing this optimal personalized preference index:
$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*}.$ (6)
The model $\pi_{\theta}$ is trained via standard supervised fine-tuning (SFT) to predict the original response $y$ given the augmented prompt $x^{\prime}$, using a cross-entropy loss:
@@ 3.3 Offline Training: Learning to Map Symbolic Tokens to Personalized Styles
$\mathcal{L}_{\text{offline}}(\theta)=-\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{t=1}^{|y|}\log\pi_{\theta}(y_{t}|x^{\prime},y_{<t})\right]$ (7)
Crucially, the LLM learns an implicit mapping: it associates the discrete token (e.g., <max_utility_index> 5) with the linguistic characteristics of responses optimal for the underlying personalized utility profile (e.g., $\hat{g}_{\psi_{5}}$). This avoids reliance on numerical interpretation, enabling more reliable personalization by learning a policy $\pi_{\theta}(y|x,i^{*})$ conditioned on the latent user personalization structure $i^{*}$.
3.4 Online Fine-Tuning: Refining Personalized Response Generation
Online fine-tuning allows the LLM to actively generate responses and iteratively refine its ability to align with diverse personalized preference profiles. Given an input prompt $x$, the model $\pi_{\theta}$ generates a response $y$. We compute its reward vector $\mathbf{z}(y)$ and evaluate it with all utility functions ${\hat{g}_{\psi_{i}}}$. The optimal utility index $i^{*}$ is determined via percentile ranking (Equation (5) relative to the current online buffer), identifying the preference profile the response best satisfies.
We further form an augmented prompt $x^{\prime}$ using $i^{*}$ as in Equation (6). High-quality pairs through rejection sampling are added to an online buffer and used to update $\pi_{\theta}$ via Equation (7). This online loop enables the model to explore beyond the initial dataset and improve its capability to generate outputs tailored to a wide range of personalized utilities.
@@ 3.5 Inference Stage
At inference time, UC-MOA offers flexible mechanisms for users to articulate their personalized multi-objective preferences, ensuring the LLM generates responses tailored to their specific needs. The user can specify their preferences in one of two ways:
Method 1: Direct Preference Vector Specification. Similar to the previous description, the user can provide a direct preference vector $\mathbf{w}=[w_{1},\ldots,w_{K}]$, where each $w_{j}$ indicates the relative importance or desired level for the $j$-th reward dimension. This vector $\mathbf{w}$ is transformed into a target reward vector $\mathbf{z}_{\text{target}}$ (e.g., via linear mapping normalized by observed reward ranges from the training data). This $\mathbf{z}_{\text{target}}$ represents the user’s ideal outcome in the multi-objective reward space. We then evaluate $\mathbf{z}_{\text{target}}$ using all pre-trained, strictly increasing utility functions $\{g_{\psi_{j}}\}_{j=1}^{M}$ from our library. The index $i^{*}$ of the utility function that best aligns with this explicit preference is selected by finding the function that maximizes the utility for $\mathbf{z}_{\text{target}}$:
$i^{*}=\arg\max_{j\in\{1,\dots,M\}}g_{\psi_{j}}(\mathbf{z}_{\text{target}}).$ (8)
This $i^{*}$ identifies the pre-trained personalized utility profile that most closely matches the user’s explicitly stated preference trade-offs.
Method 2: Preference Elicitation from Labeled Examples. Alternatively, if a user finds it more intuitive to demonstrate their preferences through examples rather than specifying a numerical vector, they can provide a small set of $L$ labeled examples, $\mathcal{D}_{\text{user}}=\{(x_{l},y_{l},s_{l})\}_{l=1}^{L}$. Each example consists of a prompt $x_{l}$, a corresponding response $y_{l}$ (which could be generated by a baseline model or written by the user), and a scalar preference score $s_{l}$ assigned by the user to that response, reflecting how well $y_{l}$ meets their personalized criteria for that specific prompt.
For each example $(x_{l},y_{l},s_{l})$ in $\mathcal{D}_{\text{user}}$:
  1. 1.
    We first compute the multi-dimensional normalized reward vector $\mathbf{z}(x_{l},y_{l})\in[0,1]^{K}$ for the response $y_{l}$.
  2. 2.
    Then, for each pre-trained utility function $\hat{g}_{\psi_{j}}$ in our library, we calculate its output score $\hat{g}_{\psi_{j}}(\mathbf{z}(x_{l},y_{l}))$.
To identify the utility function from our library that best represents the user’s demonstrated personalized preferences, we select the index $i^{*}$ that minimizes the Mean Squared Error (MSE) between its outputs and the user’s provided scores across all labeled examples:
$i^{*}=\arg\min_{j\in\{1,\dots,M\}}\frac{1}{L}\sum_{l=1}^{L}\left(g_{\psi_{j}}(\mathbf{z}(x_{l},y_{l}))-s_{l}\right)^{2}.$ (9)
This $i^{*}$ corresponds to the pre-trained personalized utility profile whose valuation of responses most closely mimics the user’s scoring behavior on their provided examples. This method allows for personalization even when users cannot easily articulate their preferences as a numerical vector.
Generating the Personalized Response. Once the optimal utility index $i^{*}$ has been determined (either via Equation (8) or Equation (9)), it signifies the personalized preference structure the LLM should adopt. The original input prompt $x_{\text{new}}$ (for which the user desires a personalized response) is augmented using the symbolic token format:
$x^{\prime}_{\text{new}}=\text{\#\#\# Prompt: }\{x_{\text{new}}\}\;\texttt{<max\_utility\_index>}+i^{*}.$ (10)
The fine-tuned LLM $\pi_{\theta}$ then generates a response conditioned on $x^{\prime}_{\text{new}}$. This response is tailored to the user’s specified or demonstrated personalized multi-objective preferences, as captured by the selected utility function $g_{\psi_{i^{*}}}$, without requiring the LLM to directly process raw numerical reward targets or complex preference descriptions at inference time.
3.6 Theoretical Guarantees
UC-MOA’s symbolic conditioning not only avoids numerical reasoning pitfalls but also provides theoretical underpinnings for achieving personalized alignment that is distributionally Pareto-optimal. Let $\mathcal{Z}\!=\![0,1]^{K}$ be the normalized $K$-dimensional reward space and $\mathcal{U}^{\uparrow}$ represent the set of strictly increasing utility functions $g_{\psi_{i}}:\mathcal{Z}\!\rightarrow\!\mathbb{R}$. Given a prompt $x$ and a utility index $i\!\in\!\{1,\dots,M\}$, our fine-tuned model $\pi_{\theta}$ exposes a utility-conditioned sub-policy
$\pi_{\theta,i}(\,\cdot)\;=\;\pi_{\theta}\!\bigl{(}\,\cdot\mid x,\langle\texttt{max\_utility\_index}\rangle+i\bigr{)}.$
First, we have the definition of the expected utility:
Definition 1 (Expected utility).
For a policy $\pi$, let $\mu_{\pi}$ denote the distribution of return vectors. Then we can define the expected utility of $\pi$ under the utility function $g_{\psi_{i}}$ as
$J_{g_{\psi_{i}}}(\pi)\;=\;\mathbb{E}_{\mathbf{z}\sim\mu_{\pi}}\bigl{[}g_{\psi_{i}}(\mathbf{z})\bigr{]}.$
Then, we hold the following assumption:
Assumption 1 (Utility-optimal sub-policies for Personalization).
For every $i\!\in\!\{1,\dots,M\}$, $\pi_{\theta,i}\in\arg\max_{\pi}J_{g_{\psi_{i}}}(\pi),$ i.e. each sub-policy maximizes the expected value of its associated strictly increasing utility $g_{\psi_{i}}$.
Finally, we present the following theorem:
Theorem 1 (Utility-Conditional Distributional Pareto-Optimality).
Under Assumption 1, for any generated response $y\!\sim\!\pi_{\theta,i}(x)$ with reward vector $\mathbf{z}(y)$, let
$i^{\star}\;=\;\arg\max_{j\in\{1,\dots,M\}}g_{\psi_{j}}\!\bigl{(}\mathbf{z}(y)\bigr{)}.$
Then the sub-policy $\pi_{\theta,i^{\star}}$ is distributionally Pareto-optimal.
A policy being distributionally Pareto-optimal means no other policy stochastically dominates it on all reward dimensions while being strictly better on at least one. Theorem 1 (proof in Appendix F) ensures that by selecting among sub-policies optimized for diverse, strictly increasing utility functions (each representing a distinct personalized preference profile), UC-MOA achieves personalized responses that are also distributionally Pareto-optimal. This result grounds UC-MOA’s symbolic token mechanism in a rigorous optimality guarantee: every utility index corresponds to a distinct, nondominated, and distributional trade-off in the multi-objective reward space.
@@ 4 Evaluation
This section details the empirical evaluation of our proposed method. We first describe the experimental setup, including the tasks, baselines, and implementation details. Subsequently, we present a comprehensive assessment of our method’s effectiveness. This is primarily demonstrated through a user study designed to capture qualitative preference alignment, and further substantiated by quantitative metrics focusing on inference consistency and Pareto front optimality . Finally, we analyze the computational efficiency of our approach. More experimental results can be found in Appendix E. We provide our code in the anonymous link https://anonymous.4open.science/r/UC-MOA-8D36.
@@ 4.1 Experiment Setup
Tasks. Following RiC [13], we select two text generation tasks for our experiments, i.e., Helpful Assistant task [2] and the Reddit Summary task [3]. For the Helpful Assistant task, we utilize three open-sourced reward models on Huggingface to measure three different objectives “harmless”, “helpful”, and “humor”. We focus on three multi-objective subtasks: (1) “harmless” vs “helpful”; (2) “humor” vs “helpful”; (3) “harmless” vs “humor”. Regarding the Reddit Summary task, we consider three reward models namely “summary”, “hallucinated”, and “faithful” to evaluate different perspectives of the summary quality. We focus on three multi-objective subtasks: (1) “summary” vs “faithful”; (2) “hallucinated” vs “faithful”; (3) “summary” vs “hallucinated”. More details on datasets and reward models are in Appendix C.
@@
Baseline methods. We compare our method with three state-of-the-art multi-objective alignment methods for personalization, i.e., Multi-Objective Reinforcement Learning from Human Feedback (MORLHF), Rewarded Soups, and Reward-in-Context (RiC). We also report the performance of the base LLaMA 2 7B model and the supervised fine-tuned (SFT) model. We document the implementation of the baseline methods in Appendix D.3.
@@
Implementation. We implement the proposed method using PyTorch. We train the machine learning models on a server with 8 NVIDIA A100 80GB GPUs and 4TB memory for all methods. We select the LLaMA 2 7B-hf model as the base model for alignment following the setting of [13]. We list the implementation details of our method (e.g., network structure for the parameterized utility function, hyper-parameter choices) in Appendix D.
4.2 Evaluating Effectiveness
We evaluate the effectiveness of our method through both qualitative user assessments and a suite of quantitative metrics.
4.2.1 User Study: Qualitative Preference Alignment
Design. To directly assess how well the models satisfy genuine user preferences, we perform a single-blind A/B test on the humorhelpfulness trade-off. We first curate five diverse dialogue prompts that explicitly ask the assistant to be either funny or helpful. For each prompt, we generate two candidate answers: one with the state-of-the-art method RiC and one with our method UC-MOA. The order of the two answers is randomly shuffled and shown side-by-side to twelve independent volunteer raters, none of whom are authors and all of whom were blind to system identity. Each rater casts one vote per prompt for the response that better satisfies both objectives, yielding a total of $5\times 12=60$ binary votes. The prompts, outputs, and anonymized comments are included in Appendix G. Our user study has an IRB approval.
Results. UC-MOA was preferred in $46$ out of the $60$ comparisons (76.7%), whereas RiC was preferred in $14$ (23.3%). A two-sided binomial test rejects the null hypothesis that the two systems are equally likely to be chosen ($p<0.001$). Qualitatively, raters commented that UC-MOA answers felt more on-point and balanced comedy with solid advice, whereas RiC outputs were occasionally funny but off–topic or helpful but bland. These findings corroborate the quantitative superiority of UC-MOA and highlight its ability to internalize nuanced, non-linear user utilities.
4.2.2 Quantitative Evaluation of Effectiveness
We further assess our method’s effectiveness using several quantitative measures.